# From Light-Cone to Supersonic Propagation of Correlations by Competing Short- and Long-Range Couplings

### v1

@@ Abstract
A fundamental question attracting enormous interest in recent decades is whether there are limits on how fast physical effects can propagate in quantum systems Lieb and Robinson (1972); Chen et al. (2023); Defenu et al. (2024, 2023). Beside its fundamental character, the answer to this question is also relevant for quantum technologies, in particular, for the preparation time of sought-after states, quantum information links and for quantum operations more generally. In this context obtaining deep insights into how correlations spread on different platforms enables one to select the most suitable one based on their speed and versatility in performing desired applications.
@@
In typical quantum systems with short-range interactions the seminal work of Lieb and Robinson has shown that correlations can only travel with a velocity that is bounded from above Lieb and Robinson (1972). The Lieb-Robinson bound can be applied to various topics Chen et al. (2023), e.g. the area law of entanglement Hastings (2007); Van Acoleyen et al. (2013), the exponential clustering for correlation functions Nachtergaele and Sims (2006), or the scrambling of information Roberts and Swingle (2016). Although the original derivation of this bound is not strictly applicable to bosonic systems, an effective light cone also exists for them at low densities Takasu et al. (2020); Kuwahara and Saito (2021a); Yin and Lucas (2022) and in the Bose-Hubbard model Kollath et al. (2007); Läuchli and Kollath (2008); Cheneau et al. (2012); Barmettler et al. (2012); Carleo et al. (2014); Faupin et al. (2022a, b). Beyond light-cone propagation is typically only present in tailored models  Eisert and Gross (2009); Kuwahara et al. (2024); Wang and Hazzard (2020). The typical light-cone spreading of correlations for short-range couplings has been observed experimentally with ultra-cold atoms in optical lattices Cheneau et al. (2012); Barmettler et al. (2012).
@@
The robustness of the light-cone spreading of correlations in systems with short-range interactions opened the question whether long-range interactions lead to a different propagation of the correlations, since they directly connect distant sites Defenu et al. (2024, 2023). An important class of long-range interactions are algebraically decaying, i.e. $x^{-\alpha}$, present in trapped ions, Rydberg, and dipolar systems Defenu et al. (2023). The role of algebraically decaying interactions for correlation spreading has been experimentally investigated with trapped ions Richerme et al. (2014); Jurcevic et al. (2014). While, in general, the concept of Lieb-Robinson bounds is not applicable to arbitrary long-range interacting systems Hastings (2012); Eisert et al. (2013); Foss-Feig et al. (2015); Schneider et al. (2021); Vu et al. (2024), other approaches pointed out that for sufficiently large values of $\alpha$ Eisert et al. (2013); Hauke and Tagliacozzo (2013); Cevolani et al. (2015); Chen and Lucas (2019); Tran et al. (2020); Kuwahara and Saito (2020); Foss-Feig et al. (2015); Else et al. (2020); Chen and Lucas (2021); Tran et al. (2021); Kuwahara and Saito (2021b); Faupin et al. (2022a, b); Gong et al. (2023); Vu et al. (2024), or special setups Jünemann et al. (2013), the correlation propagation is still effectively bound.
Still, experimentally relevant examples where unbounded correlation spreading is observed are rare and the transition from light-cone like dynamics to the latter by competing short- and long-range interaction is not yet explored. For instance, an open question in this field is whether in the presence of long-range interactions, with sufficiently small $\alpha$, their effect will dominate the shorter-range processes and simply wash out all notion of locality leading to an instantaneous build up of correlations over all distances.
@@
In this work, we investigate the emergence and disappearance of the light-cone dynamics in the presence of competing short- and global-range couplings in an out-of-equilibrium setting. We show that light-cone dynamics can trigger the onset of non-causal propagation of the globally interacting dynamics, where correlations can spread across the system almost simultaneously, independent of the distance. We dub this regime supersonic, distance-independent, propagation. As in such scenarios a distance-dependent bound is absent due to the long-range interactions we investigate the dynamics of correlations by performing numerically exact simulations for systems with short- and global-range couplings, in the presence and absence of dissipation. We find that a crossover between a light-cone behavior and the supersonic spreading of correlations arises. We supplement our simulations based on matrix product states with approximate numerical and analytical techniques to understand the key ingredients necessary for the supersonic propagation. We identify that the fluctuations of the global coupling act as the carriers of the long-range correlations, however, the exact nature of these fluctuations is not important. In particular, we show that the supersonic distance-independent propagation appears for a Bose-Hubbard atomic model, in the presence of: (i) a global-range coupling to a bosonic mode; (ii) a global purely atomic interaction; (iii) a global fluctuating field. Additionally, our findings are robust against dissipation and independent from the dimensionality of the system, making the observed effects experimentally realizable, allowing one to explore the interplay of correlations spreading and non-local dissipation.
Setup: We consider one of the most prominent platforms for controlling the interplay between short- and long-range couplings, interacting bosonic atoms confined to an external optical lattice and coupled to the field of a high-finesse cavity Ritsch et al. (2013); Mivehvar et al. (2021). A lot of theoretical and experimental effort has been focused on the nature of steady states in such models Klinder et al. (2015); Landig et al. (2016); Hruby et al. (2018); Ritsch et al. (2013); Niedenzu et al. (2010); Silver et al. (2010); Fernández-Vidal et al. (2010); Li et al. (2013); Bakhtiari et al. (2015); Flottat et al. (2017); Lin et al. (2019); Himbert et al. (2019); Halati et al. (2020a); Bezvershenko et al. (2021); Sharma et al. (2022); Chanda et al. (2022), and certain aspects of their dynamics Rodríguez Chiacchio and Nunnenkamp (2018); Halati et al. (2022, 2025). The evolution of the density matrix $\hat{\rho}$ is described by the Lindblad master equation Carmichael (1991); Breuer and Petruccione (2002); Maschler et al. (2008); Ritsch et al. (2013); Mivehvar et al. (2021)
$\partialderivative{t}\hat{\rho}=-\frac{i}{\hbar}\left[\hat{H},\hat{\rho}\right]+\frac{\Gamma}{2}\left(2\hat{a}\hat{\rho}\hat{a}^{\dagger}-\hat{a}^{\dagger}\hat{a}\hat{\rho}-\hat{\rho}\hat{a}^{\dagger}\hat{a}\right),$ (1)
@@
where $\Gamma$ is the strength of the cavity losses. The Hamiltonian is given by
@@
The local processes consist in the repulsive on-site interactions of strength $U$ and atomic tunneling with the amplitude $J$ between neighboring sites $\langle j,j^{\prime}\rangle$. The effective light-matter coupling strength $\Omega$ is controllable by a transverse standing-wave laser beam, where $\delta$ is the detuning of the cavity with respect to the pump frequency. The period of the lattice is chosen to be twice the period of the cavity mode. This effectively couples the cavity to the density imbalance of the atoms of a bipartite lattice, $\hat{\Delta}=\sum_{j\in A}\hat{n}_{j}-\sum_{j\in B}\hat{n}_{j}$, with the sublattices $A$ and $B$.
Figure 1: (a), (c), (e) The space-time propagation of the correlations $\mathcal{C}_{nn}(d,t)$, computed with $\hat{H}$, Eq. (2), for the atoms-cavity coupling strength $\hbar\Omega/J\in\{0.03,0.42,0.99\}$. The blue line is a guide to the eye and approximates the front of the light-cone propagation for $\Omega=0$. (b), (d), (f) The dynamics of $\mathcal{C}_{nn}(d,t)$ for several distances and $\hbar\Omega/J\in\{0.03,0.42,0.99\}$. The curves depicted with shades of orange to brown correspond to $\hat{H}$, Eq. (2), the dashed curves of shades of green to $\hat{H}_{\text{atom-only}}$, Eq. (4), and the blue curves to the $\Omega=0$ case. The magenta lines represent algebraically scaling $\propto t^{\alpha}$, with $\alpha\in\{2,4,8\}$. The parameters used are $N=10$ particles, $L=20$ sites, $U/J=2$, $\hbar\delta/J=2$, $\hbar\Gamma/J=0$.
Quench scenario: We simulate the propagation of the correlations in a quench scenario in which initially the atoms are in an uncorrelated product state with one atom every two sites, |$\ket{1010\dots}$, and the cavity is empty. The initial state breaks the $\mathbb{Z}_{2}$ symmetry of $\hat{H}$, $(\hat{a},\hat{\Delta})\!\to\!(-\hat{a},-\hat{\Delta})$. We analyze the dynamics of connected density-density correlations
@@
where we average over all the correlations at a certain distance $d\equiv|j-j^{\prime}|$ with $\mathcal{N}$ being the number of sites at the given distance $d$. In $\mathcal{C}_{nn}(d,t)$ we subtract the disconnected part of the density-density correlations, $\left\langle\hat{n}_{j}\right\rangle\left\langle\hat{n}_{j^{\prime}}\right\rangle$, which describes the reorganization of the atomic density in the cavity-induced potential.
@@
As in the initial state the different sites are uncorrelated, this scenario allows us to investigate cleanly the space-time propagation of correlations through the system. Let us remark that one particularity of the considered correlations is that in the limit of vanishing tunneling, $J\!\!=0$, $\hat{H}$ commutes with density operators and the correlations $\mathcal{C}_{nn}$ would be constant in time. We investigate the propagation of correlations for $J\!\neq\!0$, in the presence of competing short-range and global-range terms for the 1D version of $\hat{H}$, Eq. (2), and of the Liouvillian, Eq. (1), using a recently developed method based on time-dependent matrix product state (tMPS) techniques employing swap gates for the cavity coupling and the quantum trajectories approach for the dissipative dynamics Halati et al. (2020b); supp .
@@
Light-cone to supersonic evolution: We now investigate the crossover between the light-cone evolution and distance-independent propagation of the correlations in Fig. 1, for the 1D chain of atoms coupled to the optical cavity, assuming no cavity losses ($\Gamma\!=\!0$). For low coupling to the cavity field [Fig. 1(a) for $\hbar\Omega/J=0.03$] and the considered distances a light-cone propagation is found following the blue line, which marks approximately the speed limit for the propagation of the correlations we found in the absence of the global-range coupling. For the same value of the coupling, we present in Fig. 1(b) $\mathcal{C}_{nn}(d,t)$ for certain distances as a function of time. The correlations exhibit an algebraic increase with time until reaching a maximum following the light-cone. By comparing with the $\Omega=0$ case we observe for correlations at longer distances, $d\gtrsim 10$, finite values outside of the light-cone. This can be understood by expanding the time-evolution operator, i.e. $e^{-i\hat{H}t/\hbar}\!\approx\!e^{-i\hat{H}_{\text{BH}}t}+O(t^{2}\Omega)$. Thus, the time-evolution is dominated by the Bose-Hubbard evolution beside corrections on the order of $O(t^{2}\Omega)$, or higher depending on the initial state and the form of the coupling supp .
@@
The competing behavior becomes much more apparent as we increase the strength of the coupling to the cavity, $\hbar\Omega/J=0.42$ in Fig. 1(c)-(d). In Fig. 1(c) we observe a coexistence of the light-cone induced by short-range interaction and a distance-independent increase of correlations, for $d>4$, due to the presence of the cavity-induced global-range interactions. To be more precise for the time-evolution of the correlations at different distances in Fig. 1(d) we have an initial increase of $\mathcal{C}_{nn}(d,t)$ for $d\leq 3$ one after the other, while for $d\geq 5$ all correlations increase simultaneously. The fact that we observe at very short distances and times $tJ/\hbar\lesssim 1$ the emergence of correlations following the light-cone is due to our choice of the initial state and observables. The density correlations commute with the entire Hamiltonian except the tunneling terms, implying that initially a change in the correlations can only be induced by tunneling. Thus, the cavity-mediated long-range interactions can start spreading the correlations only triggered by the short-range hopping term. By increasing the coupling strength to the cavity even further, $\hbar\Omega/J\!=\!0.99$ in Fig. 1(e)-(f), we observe that the instantaneous spreading of the correlations occurs at a shorter time, compared to the parameters in Fig. 1(c)-(d), and only for $d=1$ and $d=2$ the correlations start growing beforehand. Furthermore, for such strong long-range interactions we do not observe any remnant of the light-cone dynamics, the role of the short-range processes being restricted to generating locally the density-density correlations, which propagate via the cavity coupling.
@@
Figure 2: (a)-(b) The dependence of the correlations $\mathcal{C}_{nn}$ as a function of rescaled time $t\sqrt{\Omega J/\hbar}$ for (a) $d=5$ and (b) $d=16$, obtained for the atom-cavity Hamiltonian $\hat{H}$, Eq. (2). (c)-(d) The dependence of the correlations $\mathcal{C}_{nn}$ as a function of rescaled time $t\Omega^{2/3}(J/\hbar)^{1/3}$ for (a) $d=5$ and (b) $d=16$, obtained for the atom-only Hamiltonian $\hat{H}_{\text{atom-only}}$, Eq. (4). The magenta lines represent the algebraic scaling (a)-(b) $\propto t^{8}\Omega^{4}$ and (a)-(b) $\propto t^{6}\Omega^{4}$. The values of the coupling are $0.03\leq\hbar\Omega/J\leq 1.02$, and the same parameters as in Fig. 1.
@@
The minimal ingredients necessary to understand the distance-independent spreading are a simplified tunneling term and the global coupling to the cavity supp . Considering $\hbar\Omega\gg J$ the correlations build up following the scaling $\mathcal{C}_{nn}(d,t)\!\propto\!\Omega^{4}J^{4}t^{8}\left[\langle(\hat{a}+\hat{a}^{\dagger})^{2}\rangle-\langle\hat{a}+\hat{a}^{\dagger}\rangle^{2}\right]$, if the initial state is an uncorrelated Fock state, where the contribution is proportional to the fluctuations present in the cavity field. In Fig. 2(a),(b) we show that this scaling is dominant in the large coupling regime taking the atoms-cavity model $\hat{H}$ by collapsing the numerical results for different values of the coupling $\Omega$ onto a single curve using $(t\sqrt{\Omega J})^{8}$. Let us note that this scaling changes when other energy scales, such as the full kinetic term, the detuning $\delta$, or preexisting long-range density correlations in the initial state, become relevant.
Generalization to atom-only global interactions: In order to be able to make a more general statement regarding the phenomenology of the spreading of correlations in systems with both short-range and global-range interactions, we contrast the results of the coupled atoms-cavity system with the following atom-only model,
@@
This model represents an atom-only description of the complex hybrid system obtained by eliminating the cavity field in the limit $J,U,\hbar\Omega^{2}/\delta\ll\hbar\delta$ Ritsch et al. (2013); Mivehvar et al. (2021), and can include global range dissipative processes Fernández-Vidal et al. (2010); Mivehvar et al. (2021); Rodríguez Chiacchio and Nunnenkamp (2018); Jäger et al. (2022). $\hat{H}_{\text{atom-only}}$ includes directly the effective global-range nature induced by the coupling to the cavity in $\hat{H}_{\text{glo}}$. In the following, we complement the results obtained with the full atoms-cavity Hamiltonian, $\hat{H}$, with results obtained with $\hat{H}_{\text{atom-only}}$. The simulations of the atom-only model, Eq. (4), make use of the matrix product states implementation of the time-dependent variational principal (TDVP) Haegeman et al. (2011, 2016); supp . We show that even if the global range coupling has a different form, stemming from purely atomic global interactions, the same crossover in the correlation spreading occurs as for the coupling to the cavity field.
@@
In Fig. 1(b),(d), showing the dynamics of $\mathcal{C}_{nn}(d,t)$ for several distances, we plot both the results for the atoms-cavity Hamiltonian, Eq. (2), (continuous lines), and the results for the corresponding atom-only model, Eq. (4), (dashed lines). At small values of the long-range interactions, Fig. 1(b), we obtain a very good quantitative agreement in the dynamics of the density-density correlations. In Fig. 1(d) we observe that the initial rise of the correlations $d=1$ and $d=2$ agrees in the two models, confirming that they are mostly influenced by the short-range terms. While in both cases we have a simultaneous increase of $\mathcal{C}_{nn}(d>3,t)$, the rise of $\mathcal{C}_{nn}(d>3,t)$ for the atom-only model occurs at slightly earlier times than for the atoms-cavity model. In particular, the scaling we find is following more closely $\mathcal{C}_{nn}(d,t)\!\propto\!\Omega^{4}t^{6}$, as shown in Fig. 2(c),(d). We note that by decreasing the photonic time-scale, i.e. increasing $\delta$, while keeping $\Omega^{2}/\delta$ constant, the difference between the long distance correlations of the two models decreases (not shown) and eventually they agree in the limit in which $\hbar\delta$ is the largest energy scale. In this regime also the atoms-cavity model shows the scaling $\mathcal{C}_{nn}(d,t)\!\propto\!\Omega^{4}t^{6}$, implying that the main effect of the fast photons is to introduce the effective global interaction. The qualitative agreement in the behavior of the correlations in the two models in this regime shows that the main features of the competition of the light-cone and the distance-independent propagation due to the global range interactions are more generic than the case of hybrid atoms-cavity systems.

### v2

@@ Abstract
In recent decades, a fundamental question attracting enormous interest is what are the speed limits on how fast physical effects can propagate in non-relativistic quantum systems Lieb and Robinson (1972); Chen et al. (2023); Defenu et al. (2024, 2023). Beside its fundamental character, this problem is also relevant for quantum technologies, as it bounds the timescale of quantum operations and algorithms Deffner and Campbell (2017). The knowledge into how correlations propagate enables one to select the most suitable platform to engineer the dynamical features leading to the desired speed.
@@
In quantum systems with short-range interactions Lieb and Robinson have shown that correlations can only travel within a light-cone defined by a maximal velocity Lieb and Robinson (1972). The Lieb-Robinson bound is a key concept across different topics in quantum many-body physics, quantum thermodynamics, and quantum information Chen et al. (2023), e.g. the exponential clustering for correlation functions Nachtergaele and Sims (2006), the scrambling of information Roberts and Swingle (2016), and even the area law of entanglement Hastings (2007); Van Acoleyen et al. (2013). Although the original derivation of this bound is not strictly applicable to bosonic systems, an effective light-cone can be derived at low densities Takasu et al. (2020); Kuwahara and Saito (2021a); Yin and Lucas (2022) and in the Bose-Hubbard model Kollath et al. (2007); Läuchli and Kollath (2008); Cheneau et al. (2012); Barmettler et al. (2012); Carleo et al. (2014); Faupin et al. (2022a, b), while the spreading of correlations breaking the light-cone bound is typically only present in tailored models Eisert and Gross (2009); Kuwahara et al. (2024); Wang and Hazzard (2020). The typical light-cone spreading of correlations for short-range couplings has been observed experimentally with ultra-cold atoms in optical lattices Cheneau et al. (2012); Barmettler et al. (2012).
@@
The extension of these concepts to long-range interacting systems, where the interactions can directly connect distant sites, is non-trivial and partially counterintuitive Defenu et al. (2024, 2023). An important class of long-range interactions are algebraically decaying with the distance, i.e. $x^{-\alpha}$, naturally occurring in trapped ions, Rydberg, and dipolar systems Defenu et al. (2023). The exponent can also be tailored by engineering interactions with lasers and resonators Periwal et al. (2021); Schneider et al. (2012). While, in general, the concept of Lieb-Robinson bounds is not applicable to arbitrary long-range interacting systems Hastings (2012); Eisert et al. (2013); Foss-Feig et al. (2015); Schneider et al. (2021); Vu et al. (2024), an effective bound exists for sufficiently large values of $\alpha$ Eisert et al. (2013); Hauke and Tagliacozzo (2013); Cevolani et al. (2015); Chen and Lucas (2019); Tran et al. (2020); Kuwahara and Saito (2020); Foss-Feig et al. (2015); Else et al. (2020); Chen and Lucas (2021); Tran et al. (2021); Kuwahara and Saito (2021b); Faupin et al. (2022a, b); Gong et al. (2023); Vu et al. (2024), or special setups Jünemann et al. (2013). The spreading of correlation has been experimentally measured in trapped ion chains with tailored interactions Richerme et al. (2014); Jurcevic et al. (2014), reporting the expected deviations from the light-cone propagation Foss-Feig et al. (2015); Chen and Lucas (2021); Tran et al. (2021).
For the special case of strong long-range interactions, where $\alpha$ is smaller than the spatial dimension, theoretical bounds allow for correlations propagation that can be almost instantaneous Tran et al. (2020). On the other hand, for certain initial states, local perturbations can remain frozen, exhibiting effects analogous to localization Santos et al. (2016); Celardo et al. (2016) originating from large energy gaps in the spectra of strong long-range interacting systems. Remarkably, little is known on how correlations spread in systems with competing short- and long-range interactions. While previous work shed light on the intricate thermalization behavior in the presence of long- and short-range interactions Blaß et al. (2018); Halimeh et al. (2017); Žunkovič et al. (2018); Lerose et al. (2019); Defenu et al. (2024), theoretical analyses of correlations spreading in this situation relevant for state-of-the art experiments are rare.
@@
In this work, we investigate the dynamical spreading of correlations after a quench in a system with competing global- and short-range couplings. We consider a paradigmatic model of experimental relevance, namely, the Bose-Hubbard model of cavity quantum electrodynamics Ritsch et al. (2013); Mivehvar et al. (2021), where the short-range kinetic and on-site processes compete with global density-density interactions mediated by photon scattering. We perform numerically exact simulations based on matrix product states and supplement them with approximate numerical and analytical techniques. We characterize the interplay of light-cone dynamics and of the non-causal propagation of the globally interacting dynamics, reporting a crossover between the light-cone behavior and the supersonic spreading of correlations, where correlations can spread across the system almost simultaneously, independent of the distance. We also show that the supersonic spreading can be triggered by light-cone dynamics. While the exact simulations are crucial in obtaining the details of the non-equilibrium dynamics, the approximate approaches allow us to identify the key ingredients necessary for the supersonic propagation and to pinpoint fluctuations of the global coupling which act as the carriers of the long-range correlations.
Setup: We consider bosonic atoms in a lattice, which interact with a quantum field, representing a photonic mode mediating global interactions between the atoms. The evolution of the density matrix of the coupled system, $\hat{\rho}$, is given by the Lindblad master equation Carmichael (1991); Breuer and Petruccione (2002); Maschler et al. (2008); Ritsch et al. (2013); Mivehvar et al. (2021)
[…]
@@
where we include the losses of the cavity with rate $\Gamma$. The Hamiltonian reads
@@
$\hat{H}_{\text{BH}}=\frac{U}{2}\sum_{j}\hat{n}_{j}(\hat{n}_{j}-1)-J\sum_{\langle j,j^{\prime}\rangle}(\hat{b}_{j}^{\dagger}\hat{b}_{j^{\prime}}+\text{H.c.}).$
The local processes consist in the repulsive on-site interactions of strength $U$ and atomic tunneling between neighboring sites $\langle j,j^{\prime}\rangle$ with the amplitude $J$. The model provides a very good description of the dynamics realized in experiments Klinder et al. (2015); Landig et al. (2016); Hruby et al. (2018), where the atomic transition is far detuned and spontaneous decay can be neglected. The effective light-matter coupling strength $\Omega$ is controllable by a pump laser and $\delta$ is the cavity-pump detuning. The period of the lattice is twice the period of the cavity mode coupling the cavity to the density imbalance $\hat{\Delta}=\sum_{j\in A}\hat{n}_{j}-\sum_{j\in B}\hat{n}_{j}$ of a bipartite lattice with sublattices AA and $B$ Habibian et al. (2013); Niederle et al. (2016); Landig et al. (2016). Theoretical efforts characterized mostly the nature of steady states in such models Niedenzu et al. (2010); Silver et al. (2010); Fernández-Vidal et al. (2010); Li et al. (2013); Bakhtiari et al. (2015); Flottat et al. (2017); Lin et al. (2019); Himbert et al. (2019); Halati et al. (2020a); Bezvershenko et al. (2021); Sharma et al. (2022); Chanda et al. (2022), and certain aspects of their dynamics Rodríguez Chiacchio and Nunnenkamp (2018); Halati et al. (2022, 2025).
Figure 1: (a), (c), (e) The propagation of $\mathcal{C}_{nn}(d,t)$, computed with $\hat{H}$, Eq. (B.1), for the atoms-cavity coupling strength $\hbar\Omega/J\in\{0.03,0.42,0.99\}$. The blue line is a guide to the eye, approximating the front of the light-cone propagation for $\Omega=0$. (b), (d), (f) $\mathcal{C}_{nn}(d,t)$ for several distances and $\hbar\Omega/J\in\{0.03,0.42,0.99\}$. The curves depicted with shades of orange to brown correspond to $\hat{H}$, Eq. (B.1), the dashed curves with shades of green to $\hat{H}_{\text{atom-only}}$, Eq. (5), and the dashed blue curves to the $\Omega=0$ case. The magenta lines show the algebraically scaling $\propto t^{\alpha}$, with $\alpha\in\{2,4,8\}$. The parameters used are $N=10$ particles, $L=20$ sites, $U/J=2$, $\hbar\delta/J=2$, $\hbar\Gamma/J=0$.
@@
where we average over all correlations at a certain distance $d\equiv|j-j^{\prime}|$ with $\mathcal{N}$ being the number of sites at the given distance $d$. In $\mathcal{C}_{nn}(d,t)$ we subtract the disconnected part of the density-density correlations, $\left\langle\hat{n}_{j}\right\rangle\left\langle\hat{n}_{j^{\prime}}\right\rangle$, which describes the reorganization of the atomic density in the cavity-induced potential.
@@
As in the initial state the different sites are uncorrelated, this scenario allows us to investigate cleanly the space-time propagation of correlations through the system. We remark that one particularity of the considered correlations is that in the limit of vanishing tunneling, J=0J\!=\!0, $\hat{H}$ commutes with density operators and the correlations $\mathcal{C}_{nn}$ would be constant in time. We investigate the propagation of correlations for $J\!\neq\!0$, in the presence of competing short- and global-range terms for the 1D version of $\hat{H}$, Eq. (B.1), and of the Liouvillian, Eq. (1), using a recently developed method based on time-dependent matrix product states (tMPS) employing swap gates for the cavity coupling and quantum trajectories for the dissipative dynamics Halati et al. (2020b); supp
@@
Light-cone to supersonic evolution: We investigate the crossover between the light-cone evolution and distance-independent propagation of the correlations in Fig. 1, for the 1D chain of atoms coupled to the cavity, assuming no photon losses ($\Gamma\!=\!0$). For low coupling to the cavity field [Fig. 1(a) for $\hbar\Omega/J=0.03$] and the considered distances a light-cone propagation is found following the blue line, which marks approximately the speed limit for correlations spreading we found in the absence of the global-range coupling. For the same value of the coupling, we present in Fig. 1(b) $\mathcal{C}_{nn}(d,t)$ for certain distances as a function of time. The correlations exhibit an algebraic increase with time until reaching a maximum following the light-cone. By comparing with the $\Omega=0$ case we observe for correlations at longer distances, $d\gtrsim 10$, finite values outside of the light-cone, signaling the supersonic spreading. The light-cone spreading at small times and distances can be understood by expanding the time-evolution operator, i.e. $e^{-i\hat{H}t/\hbar}\!\approx\!e^{-i\hat{H}_{\text{BH}}t}+O(t^{2}\Omega)$. Thus, the time-evolution is dominated by the Bose-Hubbard evolution beside corrections on the order of $O(t^{2}\Omega)$, or higher depending on the initial state and the form of the coupling supp .
@@
The features of the supersonic propagation become apparent for larger values of the coupling strength, $\hbar\Omega/J=0.42$ in Fig. 1(c)-(d). Here, we observe a coexistence of the light-cone induced by short-range processes and a distance-independent increase of correlations, for $d>4$, due to the presence of the cavity-induced global-range interactions. In Fig. 1(d) we observe that $\mathcal{C}_{nn}(d,t)$ follows the light-cone at very short distances and times ($tJ/\hbar\lesssim 1$, $d\leq 3$), while for $d\geq 5$ all correlations increase simultaneously. This is due to density correlations commuting with the entire Hamiltonian except the tunneling terms, implying that an initial change in the correlations can only be induced by tunneling. Thus, the cavity-mediated long-range interactions can start spreading the correlations only triggered by the short-range hopping term. By increasing the coupling strength to the cavity even further, $\hbar\Omega/J\!=\!0.99$ in Fig. 1(e)-(f), we observe that the instantaneous spreading of the correlations occurs at a shorter times and only for $d\leq 2$ the correlations grow beforehand. Furthermore, for such strong long-range interactions we do not observe the light-cone dynamics, the role of the short-range processes being restricted to generating locally the density-density correlations, which propagate via the cavity coupling. After the initial rise, an intricate oscillatory behavior is observed in space and time, stemming from the interplay of the Bose-Hubbard terms Barmettler et al. (2012), and the odd-even global ordering induced by $\hat{\Delta}$. This behavior could be related to metastability typical of global-range interacting systems Santos et al. (2016); Chávez et al. (2021); Defenu et al. (2024) and its study is left for future work.
@@
Figure 2: (a)-(b) The dependence of the correlations $\mathcal{C}_{nn}$ as a function of rescaled time $t\sqrt{\Omega J/\hbar}$ for (a) $d=5$ and (b) $d=16$, obtained for the atom-cavity Hamiltonian $\hat{H}$, Eq. (B.1). (c)-(d) The dependence of the correlations $\mathcal{C}_{nn}$ as a function of rescaled time $t\Omega^{2/3}(J/\hbar)^{1/3}$ for (c) $d=5$ and (d) $d=16$, obtained for the atom-only Hamiltonian $\hat{H}_{\text{atom-only}}$, Eq. (5). The magenta lines represent the algebraic scaling (a)-(b) $\propto t^{8}\Omega^{4}$ and (c)-(d) $\propto t^{6}\Omega^{4}$. The values of the coupling are $0.03\leq\hbar\Omega/J\leq 1.02$, and the same parameters as in Fig. 1.
Scaling of the supersonic spreading: The short-time behavior of the distance-independent spreading can be understood with the minimal model of a simplified tunneling term and the global coupling to the cavity supp . The correlations build up following the scaling
$\mathcal{C}_{nn}(d,t)\!\propto\!\Omega^{4}J^{4}t^{8}\left[\langle(\hat{a}+\hat{a}^{\dagger})^{2}\rangle-\langle\hat{a}+\hat{a}^{\dagger}\rangle^{2}\right]$ (4)
@@
for $\hbar\Omega\!\gg\!J$ and an initial uncorrelated Fock state. Here the amplitude is proportional to the cavity field fluctuations, implying that the fluctuations are crucial to induce the supersonic spreading. In Fig. 2(a),(b) we show that this scaling is dominant in the large coupling regime of the atoms-cavity model $\hat{H}$ by collapsing the numerical results for a wide range of different values of $\Omega$ onto a single curve using $(t\sqrt{\Omega J})^{8}$. This scaling changes when other energy scales, such as the full kinetic term, the detuning $\delta$, or preexisting long-range density correlations in the initial state, become relevant. The curves at lower $\Omega$ values deviate from the scaling gradually, showing a gradual change between the light-cone and supersonic propagation of the correlations. Thus, the interplay between the kinetic process and the global coupling to the cavity field is sufficient to cause the crossover, however, the global coupling fluctuations are required in this minimal model to obtain the supersonic spreading.
Propagation mediated by atom-only global interactions: In the full model, Eq. (B.1), the cavity field acts as an explicit carrier of the long-range correlations. The question arises if a global purely atomic interaction leads to similar behavior. Thus, we contrast the results of the coupled atoms-cavity system with the atom-only model,
@@
This represents an atom-only description of the complex hybrid system obtained by eliminating the cavity field in the limit $J,U,\hbar\Omega^{2}/\delta\ll\hbar\delta$ Ritsch et al. (2013); Mivehvar et al. (2021), and can include global range dissipative processes Fernández-Vidal et al. (2010); Habibian et al. (2013); Mivehvar et al. (2021); Rodríguez Chiacchio and Nunnenkamp (2018); Jäger et al. (2022). $\hat{H}_{\text{atom-only}}$ includes directly the effective global-range nature induced by the coupling to the cavity in $\hat{H}_{\text{glo}}$. The simulations of the atom-only model, Eq. (5), make use of the matrix product states implementation of the time-dependent variational principal (TDVP) Haegeman et al. (2011, 2016); supp . We show that even for a different global-range coupling the same crossover in the correlation spreading occurs as for the coupling to the cavity field.
@@
In Fig. 1(b),(d), showing the dynamics of $\mathcal{C}_{nn}(d,t)$ for several distances, we plot both the results for the atoms-cavity Hamiltonian, Eq. (B.1), (continuous lines), and the results for the corresponding atom-only model, Eq. (5), (dashed lines). At small values of the long-range interactions, Fig. 1(b), we obtain a very good quantitative agreement in the dynamics of $\mathcal{C}_{nn}(d,t)$. In Fig. 1(d) we observe that the initial rise of the correlations $d=1$ and $d=2$ agrees in the two models, confirming that they are mostly influenced by the short-range terms. While in both cases we have a simultaneous increase of $\mathcal{C}_{nn}(d>3,t)$, the rise of $\mathcal{C}_{nn}(d>3,t)$ for the atom-only model occurs at slightly earlier times than for the atoms-cavity model. In particular, the scaling we find is following more closely $\mathcal{C}_{nn}(d,t)\!\propto\!\Omega^{4}t^{6}$, as shown in Fig. 2(c),(d). We note that by decreasing the photonic time-scale, i.e. increasing $\delta$, while keeping $\Omega^{2}/\delta$ constant, the difference between the long distance correlations of the two models decreases (not shown) and eventually they agree in the limit in which $\hbar\delta$ is the largest energy scale. In this regime also the atoms-cavity model shows the scaling $\mathcal{C}_{nn}(d,t)\!\propto\!\Omega^{4}t^{6}$, implying that the main effect of the fast photons is to introduce the effective global interaction. The qualitative agreement in the behavior of the correlations in the two models shows that the main features of the competition of the light-cone and the distance-independent propagation due to the global range interactions are more generic than the case of hybrid atoms-cavity systems.